OpenStudy (anonymous):

5t^2-7t-60=0 How to solve for t. Step by Step.

OpenStudy (maheshmeghwal9):

use quadratic formula for an equation like:- \[ax^2+bx+c=0\] as; \[[-b \pm \sqrt{b^2-4ac}]/2a.\] & get 2 solutions for "x". :)

OpenStudy (anonymous):

actually its a physics question so could you plzz show the solving

OpenStudy (maheshmeghwal9):

first solution:-\[-(-7)+\sqrt{(-7)^2-4*5*-60}/2*5\]

OpenStudy (maheshmeghwal9):

\[[7+\sqrt{49+1200}]/10\]

OpenStudy (maheshmeghwal9):

u can do this now & find another solution by replacing (-) with (+). :)

OpenStudy (anonymous):

sorry man plzz do it full plzz i beg you

OpenStudy (maheshmeghwal9):

It can be possible in physics that there is a condition of t>0, so u have to avoid (-)ve answer

OpenStudy (maheshmeghwal9):

k! accept (+)ve answer

OpenStudy (anonymous):

plz do it plzz

OpenStudy (maheshmeghwal9):

 Now it is ur duty to find